OpenStudy (anonymous):

Which of these is a simplified form of the equation 8y + 4 = 6 + 2y + 1y? 5y = 2 5y = 10 11y = 2 11y = 10

OpenStudy (anonymous):

put like terms together

OpenStudy (anonymous):

so it would be 8y-2y-1y=6-4

pooja195 (pooja195):

y+4=6+2y+1y y+4=6+2y+y y+4=(2y+y)+(6)(Combine Like Terms) y+4=3y+6 y+4=3y+6

pooja195 (pooja195):

Anyways so just continue to combine the like terms

OpenStudy (study_together):

It's D because if we simplify both the sides we get 11y=10

OpenStudy (abb0t):

It's not D.

OpenStudy (abb0t):

It's B